OpenStudy (anonymous):

How would you solve the augmented matrix 2x+y-2z=7 , x-2y-5z=-1, 4X+y+z=-1

OpenStudy (anonymous):

2x+y-2z=7..(1) x-2y-5z=-1..(2) 4X+y+z=-1...(3) (1)x2 4x+2y-4z=14..(1)' (1)'+(2) 5x-9z=13...(4) (1)-(3) -2x-3z=8...(5) (5)x3 -6x-9z=24..(4)' (4)'-(5) -11x=11 x=-1 sub to (5) -2(-1)-3z=8 -3z=6 z=-2 sub to 2x+y-2z=7..(1) 2(-1)+y-2(-2)=7 y=5